# Dense 3D Displacement Estimation for Landslide Monitoring via Fusion of TLS Point Clouds and Embedded RGB Images

### v1

@@ 3.3.2 Patch matching from 2D source
To ensure high spatial coverage of the output, we employ 2D pixel matching as a complementary approach. This approach operates on the full-resolution point clouds rather than the downsampled ones. Using the available camera intrinsic and extrinsic parameters, we first perform 2D pixel matching on the selected source and target RGB images. Then, we use a NN search to lift the established 2D pixel matches to 3D point matches, similar to the method we previously proposed in Wang et al. (2025a). However, point cloud tiles may span multiple RGB images, resulting in only a subset of points being validly projected onto any single RGB image. To address this issue, we associate each point cloud tile with its top-$k$ images, where the $k$-th image is the one with the $k$-th most projected points. Formally, for a given source point cloud tile $\mathbf{P}^{t}$, it is projected onto all available source images, and the top-$k$ images are selected. The same process is applied to the corresponding target point cloud tile. We apply image matching to the source and target images using a state-of-the-art deep-learning image matching algorithm, Efficient LoFTR (Wang et al., 2024). To integrate all 2D pixel matches across image pairs, we first select matches from the image pair with the most matches, adding matches from additional pairs only if the corresponding 3D points have not already been matched. Finally, we filter the lifted 3D point matches using a maximum displacement magnitude threshold, e.g., 10 m, to remove outliers with excessive displacement magnitudes. To construct patch matches, for each source patch, we count the number of matched target points falling within each target patch. The target patch with the highest count is considered the most match for the current source patch (see Fig. 5).
@@ 3.4 Patch matching refinement
$\mathbf{M}=\mathbf{M}_{3D}\cup\mathbf{M}_{2D}=\bigcup_{i=1}^{n}\{(\mathbf{p}_{i},\mathbf{q}_{i})\},$ (2)
where $\mathbf{M}_{3D}$ and $\mathbf{M}_{2D}$ represent patch matches established from the 3D and 2D source, respectively, and $p_{i}$ and $q_{i}$ $(i=1,...,n)$ are the 3D coordinates of the $n$ points matched in this way.
Figure 6: Patch matching refinement. To check the quality of a patch match. A MSDD criteria metric is computed that comparing the average pair distance between all pairs of two epochs, without the estimation of transformation.
We observe that mismatched patch matches will result in erroneous displacement estimates, as evidenced by the presence of outliers in the final displacement vectors. To mitigate this, we propose a refinement module that evaluates the quality of patch matches and filters out those with unreliable correspondences. Inspired by the as-isometric-as-possible constraint (Kilian et al., 2007; Huang et al., 2008) in computer graphics, our refinement module enforces local geometric consistency by preserving distances between points in the source and target patches. We introduce the Mean Squared Distance Deviation (MSDD) (cf. Eq. 3) as a measure of isometry preservation. A patch match is considered high quality if its MSDD is below a threshold $\delta_{1}$, and the proportion of squared distance deviations below $\delta_{1}$ exceeds a threshold $\delta_{2}$. These two thresholds are manually fine-tuned based on the specific characteristics of the dataset being processed. Only matches satisfying these criteria are retained for subsequent fine matching.
$\mathrm{MSDD}=\frac{1}{2N}\sum_{i=0}^{N-1}\sum_{j=0}^{N-1}\left|\|\mathbf{p}_{i}-\mathbf{p}_{j}\|^{2}-\|\mathbf{q}_{i}-\mathbf{q}_{j}\|^{2}\right|$ (3)
Here, $N$ denotes the number of correspondences (with $i\neq j$). A high MSDD value indicates significant deviation in the pairwise distances between corresponding points, reflecting distortion of their relative spatial structure and suggesting a poor-quality patch match. Conversely, a low MSDD value implies near-isometric consistency of structure, typically associated with a rigid transformation. By enforcing this criterion, we retain only geometrically consistent matches, thereby enhancing the reliability of displacement estimation.
@@ 3.5 Fine matching
Formally, for the patch matches $\mathbf{M}_{j}$, we employ the Kabsch algorithm (Kabsch, 1976) to estimate an initial rigid transformation $(\mathbf{R}_{j},\mathbf{T}_{j})$. This estimation is based on point matches $\{(\mathbf{p}_{i},\mathbf{q}_{i})\}_{i=1}^{n}$ within $\mathbf{M}_{j}$. This transformation is subsequently refined using point-to-point ICP (Besl and McKay, 1992), resulting in a refined transformation $(\hat{\mathbf{R}}_{j},\hat{\mathbf{T}}_{j})$. Although point matches are already relatively dense compared to most existing methods, to generate a more dense output, we apply the refined transformation to all source points $\mathbf{p}_{i}^{0}$ within the current patch match. We compute the 3D DVF restricted to the current patch j as:
$\mathbf{V}_{j}:\mathbf{p}_{i}^{0}\mapsto(\hat{\mathbf{R}}_{j}\cdot\mathbf{p}_{i}^{0}+\hat{\mathbf{T}}_{j}-\mathbf{p}_{i}^{0}),$ (4)
where $i$ ranges through all matched points in patch $j$ and $\mathbf{V}_{j}:(x,y,z)\mapsto(\Delta X,\Delta Y,\Delta Z)$ represents the per-point displacements in the current patch match. Finally, the 3D displacement vectors from all patch matches are collected to form the final 3D DVF:
$\mathbf{V}=[\mathbf{V}_{1},...,\mathbf{V}_{j},...,\mathbf{V}_{m}],$ (5)
Optionally, if exact point-wise correspondences are required, we further extract them based on the refined transformation estimated for each matched patch pair. For each transformed source point within a patch, we search for its nearest neighbor in the corresponding target patch. To reduce mismatches due to noise or structural variations, we discard correspondences where the point-wise distance exceeds a threshold (e.g., the average scan resolution).
As described in Section 3.2, we independently estimate displacements for each of the three hierarchical partition levels, followed by an integration step that consolidates the outputs. Level i, which offers the finest partitioning with the smallest and most numerous patches, is prioritized due to its strong consistency with the local rigidity assumption. However, to address cases of insufficient spatial coverage, we complement it with levels ii and iii. Specifically, for each point, we assign the displacement from level i if a valid estimate exists; otherwise, we fall back to level ii, and then to level iii.
@@ 4.1 Study areas and datasets
The first study area is located north of the Brienz village in Switzerland (see Fig. 7(a)). This area exhibits active slope movements with displacement rates reaching several meters per year (Krähenbühl and Nänni, 2017). These movements destabilize the upper rock mass, leading to large displacements and sporadic rockfalls. The Brienz TLS dataset (Kenner et al., 2022) was acquired using a Riegl VZ-6000 TLS scanner positioned near the road, approximately 1,500 m from the slope. The dataset captures landslide dynamics over an area of approximately 700 x 1000 m2. The scans have an average scan resolution of 0.08 m. RGB images were acquired using the built-in, calibrated 5-megapixel camera of the scanner. The camera captures images deflected by the laser mirror, fully covering the TLS field of view with dozens of overlapping images. Each image has a resolution of 2560 x 1920 pixels, a field of view of 7.2 x 5.5 (vertical x horizontal), a ground sampling distance (GSD) of approximately 0.05 m, and an exposure time of 5.5 ms. While manufacturer-supplied software (e.g., RiSCAN PRO v2.19.3 in our case) allows for automatic stitching into a panorama, we retain the individual images to preserve spatial accuracy and avoid stitching artifacts. For this study, we use TLS scans and their corresponding RGB images from two epochs: February and November of 2020. All images have the same intrinsic calibration but distinct poses relative to the scanner coordinate system, with parameters exported after the undistortion step in RiSCAN PRO. Additionally, observations from multiple total station (TS) prisms within the landslide area are used for quantitative evaluation.
@@
Figure 7: Overview of study areas. (a): RGB image (2020) of the Brienz landslide area with the TLS scanner location indicated. (b): Map of Switzerland showing the locations of the two study areas; (c) RGB image (2019) of the Mattertal landslide area with the TLS scanner location. The two scanned landslide regions are highlighted using  green rounded rectangles. Base map source from ©Swisstopo.
@@
The second study area is located in the Matter Valley in Switzerland (see Fig. 7(c)). This area contains various geologically active formations, including rock glaciers and landslides, with forested terrain in the lower part and rocky terrain at higher elevations (Möller et al., 2023; Shi et al., 2025). The area is subject to ongoing slope movements that could endanger local infrastructure and settlements. The Mattertal TLS dataset was acquired using a Riegl VZ-4000 TLS scanner positioned on stable bedrock on the opposite side of the valley, with a range of approximately 2,500 m. The dataset covers an area of approximately 3,000 x 2,000 m2. The scans have an average scan resolution of 0.30 m. For this study, we use TLS scans and RGB images from two epochs: July 2019 and July 2022. Similar to the Brienz dataset, we use RGB images captured by the built-in camera. These images have the same resolution as the Brienz dataset, but they have a GSD of approximately 0.20 m and an exposure time of 11 ms. We use RiSCAN PRO to undistort the images and export the intrinsic and pose information. Additionally, observations from multiple GNSS stations within the landslide area are used for quantitative evaluation.
@@ 4.2 Data preprocessing
The preprocessing steps include point cloud filtering and georeferencing. Point cloud filtering removes non-ground points, e.g., those from regions covered by vegetation. These points do not directly reflect the actual displacements of the landslide surface. We use CloudCompare v2.13.2 for this filtering process. The next step is georeferencing, which enables comparisons with external observations, e.g., measurements from TS prisms or GNSS stations. We implement georeferencing in two different ways according to the situation of each dataset. For the Brienz dataset, the landslide area is highly active, and a georeferencing matrix is available from the dataset provider for each epoch. Therefore, we simply apply the given transformation matrices to the two epoch scans for georeferencing. For the Mattertal dataset, however, most areas remain stable between analyzed epochs. We align the target epoch scan to the source epoch scan using RiSCAN PRO. We also derive a georeferencing matrix by aligning the source epoch scan to a Swiss digital elevation model using a global registration pipeline implemented in Open3D (Zhou et al., 2018). Georeferencing is achieved by applying the georeferencing matrix to the source epoch scan and aligned target epoch scans.
@@ 4.3 Evaluation
The evaluation of the proposed method consists of three aspects: (i) Accuracy of estimated displacement vectors is quantitatively assessed by comparison with external discrete observations, either from TS prisms or GNSS stations, and with manually picked reference points spatially closest to these observations. For each external observation, we identify the corresponding displacement vector by searching for its closest point in the source point cloud. This choice reflects the point-based nature of both the external observations and our outputs, enabling a direct one-to-one correspondence that avoids potential spatial smoothing. The accuracy of manual picking aligns with the average scan resolution. While alternative strategies ((Gojcic et al., 2021; Shi et al., 2025; Wang et al., 2025a) involving spatial averaging may improve robustness to noise, they risk diluting localized displacement signals and are therefore deferred to A for supplementary analysis. (ii) Spatial coverage is quantified by calculating the ratio of valid displacement estimates to the total number of points in the source point cloud for selected regions of interest (ROIs). To prevent counting duplicated estimates, both displacement estimates and source point clouds are downsampled using a voxel size corresponding to the average scan resolution. (iii) Qualitative evaluation is performed to assess the spatial distribution and directionality of the estimated displacement vectors for consistency with established geomorphological understanding.
@@ 5 Results and discussion
In this section, we first introduce the baseline methods used for comparison in 3D displacement estimation in Section 5.1. We then present the accuracy evaluation in Section 5.2, followed by the spatial coverage evaluation in Section 5.3. The effectiveness of the proposed fusion strategy is examined in Section 5.4, along with the discussion of the impact of multi-level partitioning in Section 5.5. We further assess the contribution of the refinement module in Section 5.6. Finally, the limitations of the proposed method and potential improvements are discussed in Section 5.7.
@@ 5.1 Baseline selection
We compare our method against three existing approaches for estimating dense 3D DVFs in landslide monitoring: Piecewise ICP (Friedli and Wieser, 2016), F2S3 (Gojcic et al., 2021), and our previously proposed RGB-Guided method (Wang et al., 2025a). These methods serve as representatives of Categories II, III and IV, as introduced in Section 2.1. Categories I and V are omitted, as the former does not yield full 3D displacement vectors, and the latter is not designed for TLS point clouds. To enable full-scene landslide analysis, we adapt the Piecewise ICP and RGB-Guided methods, which are limited to single regions of interest, by incorporating the tiling process. For other settings, we follow the choices described in the respective publications.

### v2

@@ 3.3.2 Patch matching from 2D source
To ensure high spatial coverage of the output, we employ 2D pixel matching as a complementary approach. This approach operates on the full-resolution point clouds rather than the downsampled ones. Using the available camera intrinsic and extrinsic parameters, we first perform 2D pixel matching on the selected source and target RGB images. Then, we use a NN search to lift the established 2D pixel matches to 3D point matches, similar to the method we previously proposed in Wang et al. (2025a). However, point cloud tiles may span multiple RGB images, resulting in only a subset of points being validly projected onto any single RGB image. To address this issue, we associate each point cloud tile with its top-$k$ images, where the $k$-th image is the one with the $k$-th most projected points [^2] For the number of images (top-$k$) used in image matching, the selection reflects a trade-off between computational efficiency and spatial coverage. In our framework, each top-$k$ selection results in $k^{2}$ image matching operations. We found that setting $k$ as 1 provides sufficient spatial coverage while keeping the computational cost manageable. Formally, for a given source point cloud tile $\mathbf{P}^{t}$, it is projected onto all available source images, and the top-$k$ images are selected. The same process is applied to the corresponding target point cloud tile. We apply image matching to the source and target images using a state-of-the-art deep learning-based image matching algorithm, Efficient LoFTR (Wang et al., 2024). To integrate all 2D pixel matches across image pairs, we first select matches from the image pair with the most matches, adding matches from additional pairs only if the corresponding 3D points have not already been matched.
Finally, we filter the lifted 3D point matches using a maximum displacement magnitude threshold, e.g., 10 m, to remove outliers with excessive displacement magnitudes. To construct patch matches, for each source patch, we count the number of matched target points falling within each target patch. The target patch with the highest count is considered the best match for the current source patch (see Figure 5).
@@ 3.4 Patch matching refinement
$\mathbf{M}^{(l)}=\mathbf{M}^{(l)}_{3D}\cup\mathbf{M}^{(l)}_{2D}=\big\{\,(P^{(l)}_{i},\;Q^{(l)}_{m(i)})\;\big|\;i\in\mathcal{I},\;m:\mathcal{I}\!\to\!\mathcal{J}\ \text{is injective}\,\big\},$ (3)
where $\mathbf{M}_{3D}^{(l)}$ and $\mathbf{M}_{2D}^{(l)}$ denote the patch matches established from the 3D and 2D sources at hierarchy level $l$, respectively. $P^{(l)}_{i}$ and $Q^{(l)}_{m(i)}$ represent the $i$-th source patch and its corresponding target patch, respectively (see Figure 6).
Figure 6: Patch matching refinement. To check the quality of a patch match, a MADD criterion metric is computed that compares the mean pair distance between all pairs of two epochs, without the estimation of transformation.
We observe that mismatched patch matches will result in erroneous displacement estimates, as evidenced by the presence of outliers in the final displacement vectors. To mitigate this, we propose a refinement module that evaluates the quality of patch matches and filters out those with unreliable correspondences. Inspired by the as-isometric-as-possible constraint (Kilian et al., 2007; Huang et al., 2008) in computer graphics, our refinement module enforces local geometric consistency by preserving distances between points in the source and target patches. We introduce the Mean Absolute Distance Deviation (MADD) (cf. Equation 4) as a measure of isometry preservation. A patch match is considered high quality if its MADD is below a threshold $\delta_{1}$, and the proportion of absolute distance deviations below $\delta_{1}$ exceeds a threshold $\delta_{2}$. These two thresholds are manually fine-tuned based on the characteristics of the dataset being processed (see B). Only matches satisfying these criteria are retained for subsequent fine matching.
$\mathrm{MADD}=\frac{2}{N(N-1)}\sum_{0\leq i<j<N}\left|\|\mathbf{p}_{i}-\mathbf{p}_{j}\|_{2}-\|\mathbf{q}_{i}-\mathbf{q}_{j}\|_{2}\right|$ (4)
Here, $N$ denotes the number of correspondences. A high MADD value indicates significant deviation in the pairwise distances between corresponding points, reflecting distortion of their relative spatial structure and suggesting a poor-quality patch match. Conversely, a low MADD value implies near-isometric consistency of structure, typically associated with a rigid transformation. By enforcing this criterion, we retain only geometrically consistent matches, thereby enhancing the reliability of displacement estimation.
@@ 3.5 Fine matching
Formally, for the patch matches $\mathbf{M}_{j}^{(l)}$, we employ the Kabsch algorithm (Kabsch, 1976) to estimate an initial rigid transformation $(\mathbf{R}_{j},\mathbf{T}_{j})$. This estimation is based on point matches $\{(\mathbf{p}_{i},\mathbf{q}_{i})\}_{i=1}^{n}$ within $\mathbf{M}_{j}^{(l)}$. The point matches are derived from the same sources used for patch matching, thereby eliminating the need for additional feature extraction. This transformation is subsequently refined using point-to-point ICP (Besl and McKay, 1992), resulting in a refined transformation $(\hat{\mathbf{R}}_{j},\hat{\mathbf{T}}_{j})$. Although point matches are already relatively dense compared to most existing methods, to generate a denser output, we apply the refined transformation to all source points $\mathbf{p}_{i}^{0}$ within the current patch match. We compute the 3D DVF restricted to the current patch j as:
$\mathbf{V}_{j}^{(l)}:\mathbf{p}_{i}^{0}\mapsto(\hat{\mathbf{R}}_{j}\cdot\mathbf{p}_{i}^{0}+\hat{\mathbf{T}}_{j}-\mathbf{p}_{i}^{0}),$ (5)
where $i$ ranges through all matched points in patch $j$ and $\mathbf{V}_{j}^{(l)}:(x,y,z)\mapsto(\Delta X,\Delta Y,\Delta Z)$ represents the per-point displacements in the current patch match. Finally, the 3D displacement vectors from all patch matches are collected to form the 3D DVF at hierarchy level $l$:
$\mathbf{V}^{(l)}=[\mathbf{V}_{1}^{(l)},...,\mathbf{V}_{m}^{(l)}],$ (6)
Optionally, if exact point-wise correspondences are required, we further extract them based on the refined transformation estimated for each matched patch pair. For each transformed source point within a patch, we search for its nearest point in the corresponding target patch. To reduce mismatches due to noise or structural variations, we discard correspondences where the point-wise distance exceeds a threshold (e.g., the mean scan resolution).
As described in Section 3.2, we independently estimate displacements for each of the three hierarchical partition levels, followed by an integration step that consolidates the outputs. Let $\mathbf{V}=[\mathbf{V}_{1},\dots,\mathbf{V}_{m}]$ denote the integrated displacement vectors, where each $\mathbf{V}_{j}=\mathbf{V}_{j}^{(l)}$ and $l\in\{1,2,3\}$ indicates the first valid hierarchy level in descending priority (from level i to level iii). Level i, which offers the finest partitioning with the smallest and most numerous patches, is prioritized due to its strong consistency with the local rigidity assumption. To handle cases of insufficient spatial coverage, we complement it with levels ii and iii. Specifically, for each point, we assign the displacement from level i if a valid estimate exists; otherwise, we fall back to level ii, and then to level iii.
@@ 4.1 Study areas and datasets
The first study area is located north of the Brienz village in Switzerland (see Figure 7(a)). This area exhibits active slope movements with displacement rates reaching several meters per year (Krähenbühl and Nänni, 2017). These movements destabilize the upper rock mass, leading to large displacements and sporadic rockfalls. The Brienz TLS dataset (Kenner et al., 2022) was acquired using a RIEGL VZ-6000 TLS scanner positioned near the road, approximately 1,500 m from the slope. The dataset captures landslide dynamics over an area of approximately 700 $\times$ 1000 m2. The scans have a mean scan resolution of 0.08 m. RGB images were acquired using the built-in, calibrated 5-megapixel camera of the scanner. The camera captures images deflected by the laser mirror, fully covering the TLS field of view with dozens of overlapping images. Each image has a resolution of 1920 $\times$ 2560 pixels, a field of view of 5.5 $\times$ 7.2 (horizontal $\times$ vertical), a ground sampling distance (GSD) of approximately 0.05 m, and an exposure time of 5.5 ms. While manufacturer-supplied software (e.g., RiSCAN PRO v2.19.3 in our case) allows for automatic stitching into a panorama, we retain the individual images to preserve spatial accuracy and avoid stitching artifacts. For this study, we use TLS scans and their corresponding RGB images from two epochs: February and November of 2020. All images have the same intrinsic calibration but distinct poses relative to the scanner coordinate system, with parameters exported after the undistortion step in RiSCAN PRO. Additionally, observations from multiple total station (TS) prisms within the landslide area are used for quantitative evaluation.
@@
Figure 7: Overview of study areas. (a): RGB image (2020) of the Brienz landslide area with the TLS scanner location indicated. (b): Map of Switzerland showing the locations of the two study areas; (c) RGB image (2019) of the Mattertal landslide area with the TLS scanner location. The two scanned landslide regions are highlighted using  green rounded rectangles. Base map source: ©Swisstopo.
@@
The second study area is located in the Matter Valley in Switzerland (see Figure 7(c)). This area contains various geologically active formations, including rock glaciers and landslides, with forested terrain in the lower part and rocky terrain at higher elevations (Möller et al., 2023; Shi et al., 2025). The area is subject to ongoing slope movements that could endanger local infrastructure and settlements. The Mattertal TLS dataset was acquired using a RIEGL VZ-4000 TLS scanner positioned on stable bedrock on the opposite side of the valley, with a range of approximately 2,500 m. The dataset covers an area of approximately 3,000 $\times$ 2,000 m2. The scans have a mean scan resolution of 0.30 m. For this study, we use TLS scans and RGB images from two epochs: July 2019 and July 2022. Similar to the Brienz dataset, we use RGB images captured by the built-in camera. These images have the same resolution as the Brienz dataset, but they have a GSD of approximately 0.20 m and an exposure time of 11 ms. We use RiSCAN PRO to undistort the images and export the intrinsic and pose information. Additionally, observations from multiple GNSS stations within the landslide area are used for quantitative evaluation.
@@ 4.2 Data preprocessing
The preprocessing steps include point cloud filtering and georeferencing. Point cloud filtering removes non-ground points, e.g., those from regions covered by vegetation. These points do not directly reflect the actual displacements of the landslide surface. We use CloudCompare v2.13.2 for this filtering process. The next step is georeferencing, which enables comparisons with external observations, e.g., measurements from TS prisms or GNSS stations, and ensures that the point clouds from different epochs refer to the same coordinate frame. We implement georeferencing in two different approaches according to the characteristics of each dataset. For the Brienz dataset, the landslide area is highly active. Georeferencing was performed as part of the direct georeferencing and low-level data processing of the airborne laser scanning data by the data provider. For the Mattertal dataset, most areas remain stable between the analyzed epochs. We align the target epoch scan to the source epoch scan using RiSCAN PRO. The registration process includes plane patch-based registration, followed by multi-station adjustment, resulting in an RMSE of 0.07 m between the two epoch point clouds. We also derive a georeferencing matrix by aligning the source epoch scan to a Swiss digital elevation model using a global registration pipeline (i.e., RANSAC and ICP registration) implemented in Open3D (Zhou et al., 2018). Georeferencing is achieved by applying the georeferencing matrix to the source epoch scan and the aligned target epoch scan.
@@ 4.3 Evaluation
The evaluation of the proposed method consists of three aspects: (i) The accuracy of estimated displacement vectors is quantitatively assessed by comparison with external discrete observations, either from TS prisms or GNSS stations, and with manually picked points spatially closest to these observations. For each external observation, we identify the corresponding displacement vector by searching for its closest point in the source point cloud. This choice reflects the point-based nature of both the external observations and our outputs, enabling a direct one-to-one correspondence that avoids potential spatial smoothing. The accuracy of manual picking aligns with the mean scan resolution. Additionally, instead of searching for a single estimate, we use the mean estimated displacement vectors for comparison with external observations and manually picked points, as in existing studies (Gojcic et al., 2021; Shi et al., 2025; Wang et al., 2025a). (ii) The spatial coverage is quantified by calculating the ratio of valid displacement estimates to the total number of points in the source point cloud for selected regions of interest (ROIs). To prevent counting duplicated estimates, both displacement estimates and source point clouds are downsampled using a voxel size corresponding to the mean scan resolution. (iii) The qualitative evaluation is performed to assess whether the spatial distribution and directionality of the estimated displacement vectors are consistent with established geomorphological understanding.
@@ 5 Results and discussion
In this section, we first introduce the baseline methods used for comparison in 3D displacement estimation (Section 5.1). Next, we evaluate the accuracy and spatial coverage in Sections 5.2 and 5.3, respectively. The effectiveness of the proposed fusion strategy, impact of multi-level partitioning, and contribution of the refinement module are discussed in Sections 5.45.5, and 5.6, respectively. A runtime analysis and discussion of method limitations and potential improvements follow in Sections 5.7 and 5.8, respectively.
@@ 5.1 Baseline selection
We compare our method against three existing approaches for estimating dense 3D DVFs in landslide monitoring: Piecewise ICP (Friedli and Wieser, 2016), F2S3 (Gojcic et al., 2021), and our previously proposed RGB-guided method (Wang et al., 2025a). These methods represent Categories II, III, and IV from Section 2.1, covering key approaches for dense 3D DVF estimation. Categories I and V are omitted because Category I does not provide full 3D vectors, and Category V is not applicable to TLS point clouds. To enable full-scene landslide analysis, we adapt the Piecewise ICP and RGB-guided methods, which are limited to single regions of interest, by incorporating the tiling process. For other settings, we follow the choices described in the respective publications.